# PrisonBreak: Jailbreaking Large Language Models with Fewer Than Twenty-Five Targeted Bit-flips

### v1

@@ Appendix D Impact of Attack Configurations
Table 10: PrecisionHammer Results of All DDR4 modules. Modules are grouped by DRAM vendors (A-D) and categorized by Rowhammer vulnerability (RV) into High (RV $>$ 10), Mid (1 $<$ RV $\leq$ 10), and Low (RV $\leq$ 1) categories.
Module RV (ppm) Vicuna- 7b-v1.3 Vicuna- 7b-v1.5 Vicuna- 13b-v1.5 Llama2-7b Llama2-13b Llama3-8b Qwen2-1.5b Qwen2-7b A17 64.5 100% 100% 100% 100% 100% 100% 100% 100% A2 62.7 100% 100% 100% 100% 100% 100% 100% 100% A5 52.7 100% 100% 100% 100% 100% 100% 100% 100% A16 46.4 100% 100% 100% 100% 100% 100% 100% 100% A6 45.8 100% 100% 100% 100% 100% 100% 100% 100% A8 43.1 100% 100% 100% 100% 100% 100% 100% 100% A0 38.3 100% 100% 100% 100% 100% 100% 100% 100% A18 37.5 100% 100% 100% 100% 100% 100% 100% 100% A14 29.8 100% 100% 100% 100% 100% 100% 100% 100% A13 27.8 100% 100% 100% 100% 100% 100% 100% 100% A15 24.5 100% 100% 100% 100% 100% 100% 100% 100% A12 20.3 100% 100% 100% 100% 100% 100% 100% 100% A7 14.9 99.9% 99.1% 99.7% 54.4% 99.6% 99.1% 99.1% 98.7% A1 5.7 94.5% 47.6% 63.3% 0.2% 65.9% 54.3% 46.1% 44.2% A19 5.4 99.6% 93.5% 97% 19.4% 96.9% 94.7% 93.5% 93.8% D0 5.0 91.7% 35.9% 52.4% 0% 55.5% 44.8% 32.6% 34.1% D3 3.2 80.2% 8.7% 16.7% 0% 19.5% 10.2% 4.9% 6.2% D1 3.1 96.30% 59.8% 76% 0.4% 74.8% 63.2% 56.1% 53.5% A4 2.4 22.50% 0% 0% 0% 0% 0% 0% 0% A9 2.3 64.00% 1.5% 4.4% 0% 4.4% 1.8% 0.6% 0.7% A11 1.5 38.10% 0% 0.4% 0% 0.4% 0.1% 0% 0% A10 1.4 35.20% 0% 0.2% 0% 0.1% 0% 0.1% 0% C2 1.2 23.90% 0% 0% 0% 0% 0% 0% 0% D2 0.9453 51.20% 0.2% 0.4% 0% 1% 0.3% 0% 0.1% A3 0.8130 10.70% 0% 0% 0% 0% 0% 0% 0% B9 0.5695 3.70% 0% 0% 0% 0% 0% 0% 0% B4 0.5155 1.90% 0% 0% 0% 0% 0% 0% 0% C5 0.4787 $0$% 0% 0% 0% 0% 0% C4 0.3581 $0$% 0% 0% 0% 0% 0% C3 0.2962 $0$% 0% 0% 0% 0% 0% B1 0.2356 $0$% 0% 0% 0% 0% 0% B8 0.1201 0% 0% 0% 0% 0% 0% 0% 0% B3 0.0517 0% 0% 0% 0% 0% 0% 0% 0% B6 0.0363 0% 0% 0% 0% 0% 0% 0% 0% B7 0.0326 0% 0% 0% 0% 0% 0% 0% 0% B0 0.0293 0% 0% 0% 0% 0% 0% 0% 0% C1 0.0130 0% 0% 0% 0% 0% 0% 0% 0% C0 0.0121 0% 0% 0% 0% 0% 0% 0% 0% B2 0.0070 0% 0% 0% 0% 0% 0% 0% 0% B5 0.0065 0% 0% 0% 0% 0% 0% 0% 0%

### v2

@@ Appendix D Results on tinyBenchmarks
TABLE X: Rowhammer results for all DDR4 modules. Modules are grouped by DRAM vendors (A-D) and categorized by Rowhammer vulnerability (RV) (ppm) into High (RV $>$ 10), Mid (1 $<$ RV $\leq$ 10), and Low (RV $\leq$ 1) categories. We report the Rowhammer success rate (RSR): the ratio of successful Rowhammer simulation trials (see §7.1 for the complete definition).
Module RV (ppm) Vicuna- 7b-v1.3 Vicuna- 7b-v1.5 Vicuna- 13b-v1.5 Llama2-7b Llama2-13b Llama3-8b Llama3-70b Tulu3-8b Qwen2-1.5b Qwen2-7b A17 64.5 100% 100% 100% 100% 100% 100% 100% 100% 100% 100% A2 62.7 100% 100% 100% 100% 100% 100% 100% 100% 100% 100% A5 52.7 100% 100% 100% 100% 100% 100% 100% 100% 100% 100% A16 46.4 100% 100% 100% 100% 100% 100% 100% 100% 100% 100% A6 45.8 100% 100% 100% 100% 100% 100% 100% 100% 100% 100% A8 43.1 100% 100% 100% 100% 100% 100% 100% 100% 100% 100% A0 38.3 100% 100% 100% 100% 100% 100% 100% 100% 100% 100% A18 37.5 100% 100% 100% 100% 100% 100% 100% 100% 100% 100% A14 29.8 100% 100% 100% 100% 100% 100% 100% 100% 100% 100% A13 27.8 100% 100% 100% 100% 100% 100% 100% 100% 100% 100% A15 24.5 100% 100% 100% 100% 100% 100% 100% 100% 100% 100% A12 20.3 100% 100% 100% 100% 100% 100% 100% 100% 100% 100% A7 14.9 99.9% 99.1% 99.7% 54.4% 99.6% 99.1% 99.1% 99.3% 99.1% 98.7% A1 5.7 94.5% 47.6% 63.3% 0.2% 65.9% 54.3% 57.3% 54.8% 46.1% 44.2% A19 5.4 99.6% 93.5% 97% 19.4% 96.9% 94.7% 95.1% 95.4% 93.5% 93.8% D0 5.0 91.7% 35.9% 52.4% 0% 55.5% 44.8% 46.8% 39.9% 32.6% 34.1% D3 3.2 80.2% 8.7% 16.7% 0% 19.5% 10.2% 15.2% 11.2% 4.9% 6.2% D1 3.1 96.30% 59.8% 76% 0.4% 74.8% 63.2% 68.2% 66.3% 56.1% 53.5% A4 2.4 22.50% 0% 0% 0% 0% 0% 0% 0% 0% 0% A9 2.3 64.00% 1.5% 4.4% 0% 4.4% 1.8% 3.1% 2.5% 0.6% 0.7% A11 1.5 38.10% 0% 0.4% 0% 0.4% 0.1% 0.1% 0% 0% 0% A10 1.4 35.20% 0% 0.2% 0% 0.1% 0% 0.4% 0% 0.1% 0% C2 1.2 23.90% 0% 0% 0% 0% 0% 0% 0% 0% 0% D2 0.9453 51.20% 0.2% 0.4% 0% 1% 0.3% 0.6% 0.5% 0% 0.1% A3 0.8130 10.70% 0% 0% 0% 0% 0% 0% 0% 0% 0% B9 0.5695 3.70% 0% 0% 0% 0% 0% 0% 0% 0% 0% B4 0.5155 1.90% 0% 0% 0% 0% 0% 0% 0% 0% 0% C5 0.4787 $0$% 0% 0% 0% 0% 0% 0% 0% C4 0.3581 $0$% 0% 0% 0% 0% 0% 0% 0% C3 0.2962 $0$% 0% 0% 0% 0% 0% 0% 0% B1 0.2356 $0$% 0% 0% 0% 0% 0% 0% 0% B8 0.1201 0% 0% 0% 0% 0% 0% 0% 0% 0% 0% B3 0.0517 0% 0% 0% 0% 0% 0% 0% 0% 0% 0% B6 0.0363 0% 0% 0% 0% 0% 0% 0% 0% 0% 0% B7 0.0326 0% 0% 0% 0% 0% 0% 0% 0% 0% 0% B0 0.0293 0% 0% 0% 0% 0% 0% 0% 0% 0% 0% C1 0.0130 0% 0% 0% 0% 0% 0% 0% 0% 0% 0% C0 0.0121 0% 0% 0% 0% 0% 0% 0% 0% 0% 0% B2 0.0070 0% 0% 0% 0% 0% 0% 0% 0% 0% 0% B5 0.0065 0% 0% 0% 0% 0% 0% 0% 0% 0% 0%
Appendix E Detailed Discussion on Countermeasures
In this section, we complement §9 and provide a detailed discussion on potential countermeasures to PrisonBreak.
Rowhammer defenses. Popular Rowhammer mitigation strategies include hardware-level defenses that employ row-refreshing and swapping [82, 83, 84, 16, 85] or error correcting codes (ECC) [86, 87]. However, row-refreshing and swapping incur significant area and latency costs, and ECC are bypassable with advanced Rowhammer attacks [62, 69, 41]. In contrast, system-level defenses modify memory allocation protocols [88, 89, 90, 70], but require changes to memory allocators and specialized memory layouts, which introduce substantial memory and computational overhead. The limitations and overhead have restricted their adoption in real-world infrastructure.
Jailbreaking defenses attempt to prevent harmful generations at the prompt and model level. Prompt-based defenses include detecting anomalies in the input space [91] or modifying the system prompt to improve resilience [71, 92, 80]. PrisonBreak is naturally resilient to defenses that detect input-space changes, and we demonstrate in §9 that system prompts can be bypassed by incorporating them during search. Contrarily, model-level jailbreaking defenses focus on the model’s representations and outputs. Output-oriented works propose using separate models to detect or steer harmful outputs [72, 74] or leveraging multiple forward passes to refine responses [80, 73]. These defenses are effective but incur substantial computational costs as they require several forward passes per generated token. Additional works apply defenses before the output, analyzing gradients to detect jailbreaks [93, 94]—as our behavioral analysis finds PrisonBreak to be stealthier than prompt-based methods, we expect increased resilience.
Model-based bit-flip defenses mitigate the impact of bit-wise corruptions in parameters. They generally fall into two categories: detection and prevention. Detection defenses focus on identifying changes to weight values [95, 76, 75, 96] so that bit-wise errors can be removed by reloading the target model. However, these defenses add high overhead by defending the entire model [96, 76] or only monitor parameters critical to untargeted classification [75, 95]. A separate line of work prevents changes to weight values [77, 97, 57, 79, 78]. One approach is to use memory representations that limit changes in weight values under bit-wise corruption, such as quantization/binarization of model parameters [97, 57]. Recent work [77] instead proposes an architectural modification that introduces internal classifiers, allowing the model to stop propagating abnormal activations caused by bit-flips. While shown effective, these approaches require re-training, which incurs substantial computational costs, especially for large-scale language models. Moreover, adapting architectural modifications for computer vision models to Transformers requires non-trivial effort [98, 99].